OpenStudy (anonymous):

Solve the system using the elimination method. y-3x=2 and y=(x+1)^2-5

OpenStudy (anonymous):

The final solution is the set of all solutions. x=-2,y=-4 x=3,y=11

OpenStudy (anonymous):

wait how did you get that?

OpenStudy (anonymous):

Multiply each term in the first group by each term in the second group using the FOIL method. FOIL stands for First Outer Inner Last, and is a method of multiplying two binomials. First, multiply the first two terms in each binomial group. Next, multiply the outer terms in each group, followed by the inner terms. Finally, multiply the last two terms in each group.

OpenStudy (anonymous):

The complete solution is the set of the individual solutions. y=3x+2 x=-2,3 Multiply 3 by each term inside the parentheses. For x=-2 y=-6+2 Add 2 to -6 to get -4. For x=-2 y=-4 Multiply 3 by each term inside the parentheses. For x=3 y=9+2 Add 2 to 9 to get 11. For x=3 y=11 hope this makes a better understanding

jhonyy9 (jhonyy9):

@toomuch2handle - so than you read the text of this exercise you see that solve by elimination method - where you have doit it so ?

jhonyy9 (jhonyy9):

@happilyfalling hope will made it understandably sure - do you know what mean this ,,elimination" method ? to solving quadratic equation ?

jhonyy9 (jhonyy9):

so there are y-3x=2 y=(x+1)^2 -5 the first will be y= 3x+2 - yes ? and now multiplied by -1 both sides will get -y = -3x-2 y = (x+1)^2 -5 so now assume term by term -------------- 0 = -3x+(x+1)^2 -7 -3x + x^2 +2x +1 -7 =0 x^2 -x -6=0 do you understand it will this point please ?

jhonyy9 (jhonyy9):

not will wan being till - sorry

jhonyy9 (jhonyy9):

so than yes - you can factorizing completly the x^2 -x -6= 0 (x-3)(x+2)=0 yes ?

jhonyy9 (jhonyy9):

so this mean x-3=0 x=3 and x+2=0 x=-2

jhonyy9 (jhonyy9):

so now you return to first equation y-3x=2 and substitute in place of x theses two value of x - the 3 and -2 y-3(3)=2 y-9=2 y=11 y-3(-2)=2 y+6=2 y=-4 hope you understand it sure right easy yes ?

OpenStudy (anonymous):

thank you the both of you and sorry i never answered I was cooking.